How to solve by Elimination
2y=5x-10
3x+18=6y

Answer by oscargut(2103)

2y=5x-10
3x+18=6y

5x-2y = 10
3x-6y = -18
Multiply the first eq by (-3)
-15x+6y=-30
3x-6y = -18
adding
-12x=-48
x = 4
y = 5
I can help you with more questions at: mthman@gmail.com
Thanks